How much heat energy in kilojoules is required to raise the temperature of 456 g of copper from 43.0 ∘C to 75.8 ∘C?
1
Expert's answer
2018-01-22T05:55:08-0500
Solution
E = Cm(T2-T1), where C – copper heat capacity (0,385 kJ/kg×K)
E = 0,385×0.456×(75.8-43) = 5,76 (kJ)
Answer 5,76 kJ of heat energy are required to raise the temperature of 456 g of copper from 43.0 ∘C to 75.8 ∘C.